OpenStudy (anonymous):

Is square root 1 minus cosine squared theta = −sin Θ true? If so, in which quadrants does angle Θ terminate?

OpenStudy (anonymous):

\[\sqrt{1-\cos ^2}\theta=-\sin \theta\]

OpenStudy (anonymous):

That's how it actually looks

OpenStudy (anonymous):

@phi

OpenStudy (phi):

First, when people write the square root with no sign out front, we assume it's the positive square root. Second, there is this *very useful* identity \[ \sin^2(x) + \cos^2(x) = 1 \] from which we can derive \[ \sin^2(x) = 1 - \cos^2(x) \]

OpenStudy (anonymous):

I follow so far

OpenStudy (phi):

using that identity in your \[ \sqrt{1-\cos ^2}\theta=-\sin \theta \\ \sqrt{\sin ^2}\theta=-\sin \theta \\ \sin \theta=-\sin \theta \]

OpenStudy (anonymous):

How can something equal itself but negative?

OpenStudy (anonymous):

Ohhh, So the equation is false?

OpenStudy (phi):

when you see x = -x to solve, we add +x to both sides to get x+x= -x + x 2x = 0 divide both sides by 2 x= 0 so the solution would be x=0 in your case sin theta = 0

OpenStudy (anonymous):

Excuse me, expression

OpenStudy (phi):

it's an equation

OpenStudy (anonymous):

Darn it, I was just working with an expression, and I was thinking of that. But either way, does that mean the equation is false? Because undefined is not one of the answers.

OpenStudy (phi):

theta = 0 definitely works. I am thinking about if there are other solutions...

OpenStudy (anonymous):

I see- Would the answer choices help?

OpenStudy (anonymous):

A. False B. True; quadrants I & IV C. True; quadrants II & III D. True; quadrants III & IV

OpenStudy (phi):

ok, here is how to think about it. see https://en.wikipedia.org/wiki/Absolute_Square#Real_numbers \[ | x | = \sqrt{x^2} \] so we should write \[ \sqrt{\sin ^2}\theta=-\sin \theta \\ | \sin \theta | = -\sin \theta \\ | \sin \theta | + \sin \theta=0 \] this will be true wherever sin theta is negative. for example, say sin(theta) = -1 then | sin(theta) | + sin(theta) | -1| + -1 +1 + -1 0

OpenStudy (phi):

so it comes down to "what quadrants is sin(x) negative" ?

OpenStudy (anonymous):

1 and 2?

OpenStudy (anonymous):

Oh! negative. 3 and 4

OpenStudy (phi):

yes, 3 and 4

OpenStudy (anonymous):

So the correct answer is D?

OpenStudy (phi):

this question was tricky!

OpenStudy (phi):

yes, D is the answer

OpenStudy (anonymous):

Thank you sosososo much for your help. There's no way I would have gotten the answer alone.

OpenStudy (phi):

You could have tackled it this way: pick an angle in the "middle" of each quadrant (that you know the sine and cosine of, and test that angle in the equation. That would lead to figuring out the correct quadrants.

OpenStudy (anonymous):

Thank you! I can see how that would have worked. I actually have one more question I'm confused with. Would you possibly mind helping me out with it? I don't think it's quite as long.

OpenStudy (phi):

did you post it?

OpenStudy (anonymous):

Yes, I did. I'll mention you